OpenStudy (anonymous):

Solve the equation by graphing: x2 – 5x – 24 = 0. A. {4, 6} B. {-2, 12} C. {-4, 6} D. {-3, 8}

OpenStudy (anonymous):

Factor it as\[ (x +a ) ( x + b)= x^2 +(a+b) x +ab \] What are the two numbers so that a+b = -5 and ab =-24? a and b should be of different signs. Work it out and your roots would be -a and -b